a gun barrel weight 300 kg and has a recoil spring of stiffness 250 Newton per mm the barrel is equal to 0.8m on firing determine i the critical recoil velocity of the gun .


1
Expert's answer
2021-02-05T04:56:10-0500

(i) The natural frequency is

𝜔 = "\sqrt \frac{𝑘 }{𝑚 }"

= "\sqrt \frac{250000}{300}"

=28.87 𝑟𝑎𝑑/𝑠 .


From conservation of energy, we have

"\frac{ 𝑚𝑣2}{2}" = "\frac{𝑘𝑥2}{2 }"

𝑣 = 𝜔𝑥 = (28.87)(0.8)

=23.096𝑚/𝑠 .